A man is 37 years old and his child's age is 8. How many years ago was the product of their ages 96
Answer by ankor@dixie-net.com(22740)

A man is 37 years old and his child's age is 8.
How many years ago was the product of their ages 96
:
Let x = no. yrs ago for this to be true
:
(37-x)(8-x) = 96
FOIL
296 - 37x - 8x + x^2 = 96
x^2 - 45x + 296 - 96 = 0
x^2 - 45x + 200 = 0
Factors to
(x-40)(x-5) = 0
x = 5 yrs is the only reasonable solution
:
see if that is true
32 * 3 - 96